# BatteryML:An Open-source platform for Machine Learning on Battery Degradation

### 04 Dec 2023 v3

@@ 1 Introduction
Lithium-ion batteries, characterized by their high energy density and prolonged cycle life, have revolutionized energy storage across sectors like electric vehicles, consumer electronics, and renewable energy solutions. However, the ubiquitous adoption of these batteries comes with inherent challenges surrounding their capacity degradation and performance stability. Continuous cycling tends to diminish their charging and discharging capacities, posing dire implications for real-world applications. For instance, ”range anxiety” becomes prevalent among electric vehicle owners, and reliability concerns arise for energy storage systems. Beyond the user experience, rapid degradation introduces broader issues, such as escalating maintenance costs, heightened resource usage, environmental strain, and potential economic inefficiencies. As such, decoding and forecasting battery performance degradation has ascended as a pivotal topic in industrial artificial intelligence.
Peeling back the layers of lithium-ion batteries reveals their intricate, non-linear electrochemical dynamics (Hu et al., 2020). Degradation, as observed as diminishing performance with increased charge-discharge iterations. branches mainly into losses in lithium-ion inventory (attributed to Solid Electrolyte Interphase(SEI) film formation, electrolyte decomposition, etc.) and active material losses (like graphite delamination and binder decomposition) (Pop et al., 2007; Dubarry et al., 2012; Sarasketa-Zabala et al., 2015). Moreover, the internal resistance and excessive electrolyte losses further contribute to the battery’s declining health. Such losses in electrolytes, in particular, can precipitate a stark capacity plunge towards a battery’s lifecycle end.
Amidst this degradation complexity, reliably predicting a battery’s Remaining Useful Life (RUL) and State of Health (SOH) becomes a herculean endeavor (Lipu et al., 2018). The significance of RUL, especially in battery management, second-hand vehicle evaluation, and more, has spurred extensive research. For instance, integrating techniques like electrochemical impedance spectroscopy with machine learning has been demonstrated to hold promise (Zhang et al., 2020; Severson et al., 2019; Attia et al., 2021; 2020). Similarly, SOH estimation has seen advances through capacity-based, Coulomb counting, impedance methods, and model-based techniques, with machine learning bringing innovative dimensions (Wang et al., 2011; Plett, 2004; Barsoukov et al., 2005; Doyle et al., 1993; He et al., 2011).
@@
Challenges
  • Heterogeneous Battery Data: Battery data exhibit considerable heterogeneity. Depending on the cathode materials, distinctions include LiCoO2(LCO), LiFePO4(LFP), and LiNiMnCoO2(NMC). The diversity in data collection devices and preprocessing pipelines results in varied formats. Even for same battery chemistry, the logged charge-discharge signals are not unified in their dimensions. For example, the capacity of the battery may be reported through the areal specific capacity, the total capacity, or even normalized capacity. Manufacturers produce data outputs in unique forms, and subsequent preprocessing can further diversify these formats, such as CALCE (Xing et al., 2013; Liu et al., 2017) in xls, PCOE Ma et al. (2019) and MATR (Severson et al., 2019; Hong et al., 2020) in .mat, HUST Ma et al. (2022) in pkl, and RWTH Li et al. (2021) in csv. Furthermore, specific charge and discharge strategies differ across datasets. As a consequence, many models are developed and trained on a limited set of datasets, hindering the establishment of a universally accepted performance comparison standard.
  • Domain Knowledge for Feature Design: For machine learning professionals, battery data’s high-dimensional and heterogeneous nature demands specific domain knowledge for crafting effective feature spaces. This intricacy is a major obstacle in leveraging cutting-edge machine learning techniques for battery performance modeling.
  • Model Applicability: Battery experts, while proficient in understanding degradation mechanisms, often face challenges in building robust machine learning models. Constructing an effective model often necessitates nuanced data processing, feature engineering, and model fine-tuning. Tools and models crafted for specific data structures might not seamlessly adapt to others, restricting their broad application. Knowing the model architecture and the training data is beneficial to understanding applicability of the model the constraints.
@@
BatteryML addresses the above challenges in a holistic manner. As an inclusive open-source platform, BatteryML simplifies every stage of the process, from data preprocessing and feature extraction to applying both traditional and avant-garde models.
  • Unified Data Representation: Recognizing the challenges of diverse battery data, BatteryML introduces a standardized data representation method. It provides comprehensive processing tools to collate and harmonize virtually all public battery datasets. With this consistent data format, a uniform evaluation criterion for assessing battery degradation becomes feasible, promoting robust comparisons and insights across diverse battery contexts.
  • Comprehensive Open-source Platform: BatteryML covers essential battery research tasks like State of Charge (SOC), State of Health (SOH), and Remaining Useful Life (RUL). It offers a holistic suite of tools encompassing data preprocessing, feature and target extraction, model training, prediction, and visualization. This integrative design allows experts from varied fields to contribute, nurturing ongoing innovation in battery research.
  • Inclusion of Modern Models: BatteryML integrates a broad spectrum of tasks and models, both conventional and state-of-the-art. The platform’s modular design ensures clear demarcation between models and data processing stages, facilitating effortless integration and refinement by machine learning experts. With a standardized data format, researchers can leverage multiple datasets simultaneously, unlocking techniques like transfer learning. This fluidity not only accelerates research but also sets the stage for the integration of more sophisticated models in the coming times.
@@
Battery Domain Tasks. Lithium-ion battery lifetime modeling has been the subject of numerous studies. A vast number of researchers have proposed both physical and semi-empirical models to capture various mechanisms, including the growth of the solid-electrolyte interphase, lithium plating, active material loss, and impedance increase (Das et al., 2019; Palacín, 2018; Woosung et al., 2020). Predictive state estimation for remaining useful life in battery management systems often hinges upon these mechanistic and semi-empirical models. Specialized diagnostic measurements, such as coulombic efficiency and impedance spectroscopy, further assist in lifetime estimation (Burns et al., 2013; Chen et al., 2001; Tröltzsch et al., 2006; Love et al., 2014). Despite their success, these chemistry or mechanism-specific models often struggle with describing full cells subjected to relevant operating conditions, like fast charging. This is due to the intricate interactions between multiple degradation modes and the thermal and mechanical variances within a cell (Waldmann et al., 2014; 2015; Bach et al., 2016; Jain et al., 2013; Aykol et al., 2016).
While semi-empirical models demand in-depth battery and chemistry domain knowledge, battery degradation is inherently multifaceted, encompassing various intricate degradation mechanisms. Parallelly, the rapid evolution of machine learning and neural networks has ushered in a plethora of data-driven methodologies, including neural networks, support vector machines, and deep learning algorithms (Severson et al., 2019; Raccuglia et al., 2016; Segler et al., 2018; Nuhic et al., 2013; Ng et al., 2020; Lu et al., 2023; Chemali et al., 2018; Hu et al., 2015; won You et al., 2016; Zhang et al., 2018; Li et al., 2020; Attia et al., 2021; Ma et al., 2022; Álvarez Antón et al., 2013; Ren et al., 2018; Khumprom & Yodo, 2019; Sahinoglu et al., 2018; Jiménez-Bermejo et al., 2018; Wu et al., 2018; Guo et al., 2015; Saha et al., 2009; Hu et al., 2016; Zhang et al., 2019). These techniques, considering factors like aging, temperature, and real-world driving conditions, are instrumental in refining battery management systems by offering precise results with minimal estimation discrepancies, as highlighted in multiple reviews (Lu et al., 2013; Hannan et al., 2017; Si et al., 2011; Han et al., 2019; Rezvanizaniani et al., 2014; Berecibar et al., 2016; Li et al., 2019; Liu et al., 2020; Meng & Li, 2019; Lipu et al., 2018; Hossain Lipu et al., 2021; Ayob et al., 2022; Rauf et al., 2022; Ramadesigan et al., 2012; Xiong et al., 2018; Santhanagopalan et al., 2006; Lucu et al., 2018; Wu et al., 2016). Their integration can considerably elevate the performance, safety, and efficiency of EV batteries, spurring the broader transition to electric vehicles. Given the diverse research efforts and varying datasets, a head-to-head comparison of these methodologies is challenging. This accentuates the urgent need for a consolidated platform that standardizes battery degradation research, further propelling the field’s progression.
Battery Early Prediction Framework. The Battery Evaluation and Early Prediction Software Package (BEEP) offers an open-source, Python-centric framework designed for the efficient handling and processing of extensive battery cycling data streams (Herring et al., 2020). Notable features of BEEP encompass file-system-oriented organization of raw cycling data, validation procedures for data authenticity, parsing and molding of data into Python objects for analysis, feature extraction from organized cycling data for machine learning input, and comprehensive demonstrations for anomaly detection and cycle life early-prediction models. Although adept for private datasets, BEEP does not seamlessly integrate with the vast array of public data. Additionally, its predominant reliance on sklearn curtails support for advanced models such as CNN and LSTM, and learning paradigms like transfer learning. This limitation constricts the potential for cultivating models that generalize across diverse battery categories and operational conditions.
3 Pipeline
Figure 1: Framework of BatteryML.
@@ 3 Pipeline
As depicted in Figure 1, the BatteryML pipeline comprises an organized sequence of functional modules, guiding users through the process of model creation and application. The initial step involves converting all incoming data into a consistent format. Following this, a configuration file is crafted to specify data locations, partitioning strategies, feature and label generation methods, as well as the associated model parameters. An elaborate sample of these settings is presented in Algorithm D.1. Once configured, users can activate the comprehensive pipeline process. Modules integral to this process encompass the Train-test data split, Label extractor, Feature extractor, Data preprocessing, and the Models module. A concise walkthrough of the pipeline’s core functions, such as pipeline initialization, model training, and results extraction, is available in Appendix D.1.
  • Train-test Split Module: Users can opt for the random split methodology to allocate data into training and test subsets based on a specified proportion. Alternatively, segmentations tailored to specific dataset divisions like MATR Severson et al. (2019) and HUST Ma et al. (2022) are available. The module also offers the flexibility for custom data set partitioning.
  • Label Extractor Module: This allows for the determination of target predictions, including metrics like RUL or SOH.
  • Feature Extractor Module: To replicate one of the most popular approaches by Severson et al. (2019), the variance model method can be directly employed for feature extraction. For those seeking to implement their proprietary feature extraction techniques, the module can be tailored to accommodate such customizations.
  • Data Preprocessing Module: Here, users can independently process both features and labels, adopting strategies like log transformations or the utilization of z-score normalization.
  • Model Module: The module encompasses a broad spectrum, bifurcating into linear models, tree-based models, and neural network-based models. To reproduce specific works, it is essential to synchronize module configurations, encompassing feature extraction and model selection parameters. For an overview of the results generated by prevalent models, users can consult the baseline.ipynb documentation.
3.1 Data
@@
The multifaceted landscape of battery data stems from the diverse data collection apparatuses and preprocessing methodologies employed by manufacturers. This diversity is accentuated by the unique output formats from cyclers, further diversified by subsequent preprocessing steps. Common formats encountered include Excel, CSV, TXT, MATLAB, and Pickle. Recognizing the challenge this poses, we introduced a consolidated data format that encompasses both fundamental information and intricate charge/discharge cycle data. This unified approach comprises:
  • Specifications Data: This segment captures intrinsic battery attributes including the anode, cathode, and electrolyte materials. It also enumerates parameters such as nominal capacity, depth of charge/discharge, and the upper and lower thresholds for voltage and current.
  • Charge/Discharge Cycle Data: This section delves into the specifics of each cycle, detailing charge/discharge protocols, capacities, voltage, current, time, temperature, and internal resistance metrics.
A comprehensive outline of this data format can be found in Appendix  D.2.
Our commitment to this unified format not only facilitates efficient data management from diverse sources but also bolsters insightful comparisons into battery performance and degradation nuances. It creates a conducive environment for deploying machine learning and sophisticated data analysis strategies, propelling advancements in battery design, operations, and maintenance. For users, this format unlocks the potential to visually articulate battery data nuances. As evident in Figure 2, one can glean insights into capacity degradation trends. Such plots typically compare cycle numbers (x-axis) against capacity or normalized capacity values (y-axis), elucidating the capacity attrition over time due to degradation phenomena.
Additionally, users can delve into various battery parameters such as the trajectory of voltage curves over successive cycles or the evolution of Coulombic efficiency, as highlighted in Figure 3. Such graphical interpretations give users a richer understanding of battery performance, degradation pathways, and overarching trends. These, in turn, illuminate pathways for refinements in battery design, operations, and maintenance.
Convert Format
Before utilizing the pipeline, it is necessary to convert the raw data format to a compatible format. Here’s an example using MATR-1 and MATR-2 (Severson et al., 2019). Replace the ’/path/to/save/raw/data’ in the command with the path of the downloaded MATR data, and the preprocess function will convert each .mat file into a unified format, subsequently saving the cells as .pkl files. You can refer to the AlgorithmD.2 in Appendix to download the data and convert it to a uniform format. For detailed data format conversions, please refer to the code under the /batteryml/preprocess path.
@@ 3.2 Feature Engineering
BatteryML boasts an array of feature extraction techniques, enabling users to tailor their analyses based on distinct experimental needs. These features bifurcate into two categories: within-cycle and between-cycle.
@@
Within-cycle Features: This category encompasses characteristics observed within individual battery cycles, such as:
@@
  • QdLinear: This involves linear interpolation of the Capacity-voltage curve during discharge, yielding the Qdlinear curve.
  • Coulombic Efficiency: An indicator of how efficiently a battery can convert its stored energy.
@@
  • Internal Resistance: A measure of the opposition to current within the battery.
@@
Between-cycle Features: These features are broader, capturing trends and patterns across multiple cycles. Examples include:
  • Log Variance of $\Delta Q_{100-10}(V)$: A classical measure proposed by  (Severson et al., 2019).
  • Capacity Decay Dynamics: Linear fitting of the capacity decay curve from the 2nd to 100th cycle.
  • Average Charging Time: Calculated over the initial five cycles.
  • Temperature Dynamics: Integrated temperature changes over time, typically between the 2nd and 100th cycles.
  • Minimal Internal Resistance: The lowest resistance observed across cycles.
Moreover, users can compute tensors for neural network training by subtracting the QdLinear curve of the 10-th cycle from those spanning the 0-th to the 99-th cycles. Customizing the specific number of cycles is also feasible. The merger of within-cycle and between-cycle features provides users with enhanced analytical flexibility.
BatteryML extends beyond these standard features, offering an infrastructure where users can forge their feature combinations. For instance, the discharge model (proposed by  (Severson et al., 2019)) focuses exclusively on discharge data, generating features such as minimum, variance, skewness, and kurtosis of $\Delta Q_{100-10}(V)$, as well as differentials in discharge capacities between selected cycles. Additionally, users can craft a voltage-capacity matrix tailored for neural networks. Here, the QV curve of the discharge chain undergoes linear interpolation to yield the QdLinear curve, which spans cycle length multiplied by interpolation dimension. Subsequently, the QdLinear curve of the 10th cycle is deducted from those ranging from the 0th to 99th cycles.
Through these feature extraction methodologies, BatteryML provides users with the freedom to tailor data analyses, ensuring they align perfectly with unique research objectives and prerequisites.
3.3 Task Definition
Lithium-ion batteries, by their very electrochemical nature, are destined for performance degradation, notably evidenced by reduced charge and discharge capacities over continual cycles. Such uncontrolled degradation presents ramifications for applications down the line: eliciting ”range anxiety” in electric vehicle owners or compromising the stability of power delivery in energy storage systems. To address these concerns, Battery Management Systems (BMS) focus on precise estimations of a battery’s SOH, SOC and RUL. Aligning with this essential need, BatteryML’s task module caters to these classic battery degradation metrics: SOH, SOC, and RUL.
As batteries endure wear and tear, their performance drops, propelling the BMS to gauge the SOH, grounded in aging mechanisms and established degradation models. Multiple definitions of SOH exist, with determinants ranging from operational duration, surges in internal resistance, to experienced capacity loss. However, a prevalent notion of SOH equates it to the proportion of the current capacity to its pristine, initial capacity. Conversely, SOC is a relative measure, comparing a battery’s present capacity against its maximum, where 100% indicates a fully charged state and 0% its complete depletion. A fresh-off-the-production battery is stamped with an SOH of 100%, which slumps to 80% at its end of life (EOL).
EOL, in battery production parlance, is the stage when the capacity, post a full charge, recedes to 80% of its designated nominal value (Li et al., 2019). The remaining charge-discharge cycle count until a battery grazes its EOL is coined as its RUL. Given the majority of publicly accessible data hails from exhaustive charge-discharge tests, BatteryML leans on the proportion of discharge capacity to its original as the SOH indicator, the comparison of the battery’s current capacity against its peak as the SOC metric, and the pending charge-discharge cycles leading to the discharge-to-nominal capacity ratio meeting the benchmark (commonly 80%) as its RUL criterion. The definition of SOH and SOC is as follows.
$\mathrm{SOC}=\frac{C_{\mathrm{curr}}}{C_{\mathrm{full}}}\times 100\%,~{}~{}\mathrm{SOH}=\frac{C_{\mathrm{full}}}{C_{\mathrm{nom}}}\times 100\%$ (1)
Here, $C_{curr}$ represents the current battery capacity, $C_{full}$ signifies its maximum when fully charged, and $C_{nom}$ alludes to the nominal capacity of a brand-new battery  (Li et al., 2019).
3.4 Model Selection and Implementation
BatteryML boasts a variety of models, drawing from well-established methodologies, and is categorized primarily under linear, tree-based, and neural network paradigms. Within the linear space, it offers options such as Linear Regression and Ridge Regression. Tree-based models encompass techniques like Random Forest and LightGBM, while the neural network segment features architectures such as the MLP (Multilayer Perceptron) and CNN (Convolutional Neural Network). To cater to the diversity of experimental contexts, users have the liberty to tweak and customize these models as per their needs. A deeper dive into model intricacies can be found in AppendixD.5.

### 25 Mar 2024 v4

@@ 1 Introduction
Lithium-ion batteries, characterized by their high energy density and prolonged cycle life, have revolutionized energy storage across sectors like electric vehicles, consumer electronics, and renewable energy solutions. However, the ubiquitous adoption of these batteries comes with inherent challenges surrounding their capacity degradation and performance stability (Edge et al., 2021). Continuous cycling tends to diminish their charging and discharging capacities, posing dire implications for real-world applications. For instance, “range anxiety” becomes prevalent among electric vehicle owners, and reliability concerns arise for energy storage systems. Beyond the user experience, rapid degradation introduces broader issues, such as escalated maintenance costs, heightened resource usage, environmental strain, and potential economic inefficiencies. As such, decoding and forecasting battery performance degradation has ascended as a pivotal topic in industrial artificial intelligence.
Peeling back the layers of lithium-ion batteries reveals their intricate, non-linear electrochemical dynamics (Hu et al., 2020). Degradation, observed as diminishing performance with increased charge-discharge iterations, branches mainly into losses in lithium-ion inventory (e.g., solid electrolyte interphase film formation and electrolyte decomposition) and active material losses (e.g., graphite delamination and binder decomposition) (Pop et al., 2007; Dubarry et al., 2012; Sarasketa-Zabala et al., 2015). Moreover, the internal resistance and excessive electrolyte losses further contribute to the battery’s declining health. Such losses in electrolytes, in particular, can precipitate a stark capacity plunge towards a battery’s lifecycle end (Edge et al., 2021).
Confronting this degradation complexity, reliably predicting a battery’s remaining useful life (RUL), state of health (SOH), and state of charge (SOC) becomes a herculean endeavor (Lipu et al., 2018). The significance of RUL, especially in battery management, second-hand vehicle evaluation, and more, has spurred extensive research. For instance, integrating techniques like electrochemical impedance spectroscopy with machine learning has been demonstrated to hold promise (Zhang et al., 2020; Severson et al., 2019; Attia et al., 2021; 2020). Similarly, SOH and SOC estimation has seen advances through capacity-based, Coulomb counting, impedance methods, and model-based techniques, with machine learning bringing innovative dimensions (Wang et al., 2011; Plett, 2004; Barsoukov et al., 2005; Doyle et al., 1993; He et al., 2011).
@@
[…]
  • Domain knowledge. On the one hand, machine learning professionals struggle to craft effective feature spaces due to the high-dimensional and heterogeneous characteristics of battery data. This intricacy presents a significant challenge in applying advanced machine learning techniques to battery performance modeling.
  • Model development. On the other hand, battery experts, while proficient in understanding degradation mechanisms, often face challenges in building robust machine learning model due to the nuanced data cleaning, feature engineering, and model fine-tuning processes. Existing tools and models that arecrafted for specific data structures might not seamlessly adapt to other scenarios.
@@
BatteryML addresses the above challenges in a holistic manner. As an inclusive open-source platform, BatteryML simplifies every stage of battery modeling, from data preprocessing and feature construction to model training and inference.
  • Unified data representation. Recognizing the challenges of diverse battery data, BatteryML introduces a standardized data representation method. It provides comprehensive processing tools to collate and harmonize virtually all public battery datasets. With this consistent data representation, a uniform evaluation criterion for assessing battery degradation becomes feasible, promoting robust comparisons and insights across diverse battery contexts.
  • Comprehensive open-source platform. BatteryML covers essential battery research tasks like State of Charge (SOC), State of Health (SOH), and Remaining Useful Life (RUL). It offers a holistic suite of tools encompassing data preprocessing, feature and target extraction, model training, prediction, and visualization. This integrative design allows experts from varied fields to contribute, nurturing ongoing innovation in battery research.
  • State-of-the-art model integration. BatteryML seamlessly integrates a wide array of models, spanning both traditional and cutting-edge techniques. The platform’s modular design ensures clear demarcation between models and data processing stages, facilitating effortless integration and refinement by machine learning experts. With a unified data representation, researchers can leverage multiple datasets simultaneously, unlocking techniques like transfer learning. This fluidity not only accelerates research but also sets the stage for the integration of more sophisticated models in the coming times.
@@
Battery modeling tasks. Lithium-ion battery lifetime modeling has been the subject of numerous studies. A vast number of researchers have proposed both physical and semi-empirical models to capture various mechanisms, including the growth of the solid-electrolyte interphase, lithium plating, active material loss, and impedance increase (Das et al., 2019; Palacín, 2018; Woosung et al., 2020). Predictive state estimation for remaining useful life in battery management systems often hinges upon these mechanistic and semi-empirical models. Specialized diagnostic measurements, such as coulombic efficiency and impedance spectroscopy, further assist in lifetime estimation (Burns et al., 2013; Chen et al., 2001; Tröltzsch et al., 2006; Love et al., 2014). Despite their success, these chemistry or mechanism-specific models often struggle to accurately characterize the battery degradation on the cell level due to the intricate interactions between multiple degradation modes and the thermal and mechanical variances within a cell (Waldmann et al., 2014; 2015; Bach et al., 2016; Jain et al., 2013; Aykol et al., 2016).
While semi-empirical approaches require in-depth battery and chemistry domain knowledge to model various intricate degradation mechanisms, the rapid evolution of machine learning provides a fully data-driven methodology, using linearmodels, support vector machines, and neural networks for accurate battery degradation modeling (Severson et al., 2019; Segler et al., 2018; Ng et al., 2020; Lu et al., 2023; Chemali et al., 2018; Zhang et al., 2018; Li et al., 2020; Attia et al., 2021; Ma et al., 2022; Ren et al., 2018; Khumprom & Yodo, 2019; Sahinoglu et al., 2018; Jiménez-Bermejo et al., 2018; Wu et al., 2018; Zhang et al., 2019). Such data-driven approaches allow seamlessly integration of different degradation factors such as electrical signals, temperature and electrochemical impedance spectroscopy data, for flexible battery modeling (Han et al., 2019; Li et al., 2019; Meng & Li, 2019; Liu et al., 2020; Hossain Lipu et al., 2021; Ayob et al., 2022; Rauf et al., 2022). On one hand, the flourishing of these data-driven approaches continues to propel advancements in battery data acquisition and the performance of machine learning models. On the other hand, due to the absence of a unified modeling framework, the problem settings and data representations vary across different models, making it challenging to achieve stable replication and comparison. This accentuates the urgent need for a consolidated platform that standardizes battery degradation research, further propelling the field’s progression.
Battery Early Prediction Framewok. The Battery Evaluation and Early Prediction Software Package (BEEP) offers an open-source, Python-centric framework designed for the efficient handling and processing of extensive battery cycling data streams (Herring et al., 2020). Notable features of BEEP encompass file-system-oriented organization of raw cycling data, validation procedures for data authenticity, linear model learning for anomaly detection and cycle life early-prediction. While BEEP positions itself as a tool designed to assist battery experts with coding skills to more efficiently conduct battery life predictions to validate design ideas, BatteryML aims to bridge the gap between the battery and machine learning communities. Through its modular design, BatteryML decouples the knowledge dependencies of the two communities, enabling battery experts to utilize the most advanced machine learning models, while also allowing machine learning professionals to more effectively optimize models for battery data. BatteryML’s data representation naturally supports the transformation of battery data into multidimensional tensors, thereby facilitating seamless integration for battery experts into existing deep learning frameworks. Moreover, BatteryML empowers machine learning professionals to explore advanced learning paradigms, such as transfer learning, on a variety of battery data.
3 BatteryML Platform
Figure 1: An overview of BatteryML.
@@ 3 BatteryML Platform
As depicted in Figure 1, the BatteryML pipeline comprises an organized sequence of functional modules, guiding users through the process of model creation and application. The initial step involves converting all incoming data into a consistent format. Following this, a configuration file is crafted to specify data locations, partitioning strategies, feature and label generation methods, as well as the associated model parameters. An elaborate sample of these settings is presented in Code D.1. Once configured, the comprehensive pipeline process allows end-to-end battery degradation modeling. Components integrated in this process encompass the train-test data split module, label extractor module, feature extractor module, data preprocessing module, and the model module. Detailed examples of the pipeline’s core functions, including pipeline initialization, model training, and result extraction, is available in Appendix D.1.
  • Train-test split module. This module determines the split strategy of the learning process. BatteryML allows users to randomly allocate battery cells into training and test subsets with respect to a given split proportion. Alternatively, BatteryML also provide standard data splits for popular datasets such as MATR Severson et al. (2019) and HUST Ma et al. (2022) to enable reproducible and comparable experiments. The module also offers high flexibility for custom data partitioning.
  • Label extractor module. This module automatically annotates the prediction target for major battery modeling tasks such as RUL and SOH prediction.
  • Feature extractor module. This module implements the popular features designed by domain experts such as the cycle-difference features for battery life prediction Severson et al. (2019). This module also enables extraction of raw electric signals and conversion to tensors for training neural networks. This module is highly extensible to support custom feature design.
  • Data preprocessing module. This module supports flexible transformation and refinement independently for both features and labels, including data normalization and augmentation techniques to boost model performance.
  • Model module. This module specifies the model structure and learning parameters. BatteryML currently supports a broad spectrum of machine learning models including linear models, tree-based models, and neural network-based models. Users can effortlessly manipulate the model behavior without the concerning the intricacies of the specialized domain knowledge.
3.1 BatteryData: A Unified Battery Data Representation
@@
The multifaceted landscape of battery data stems from the diverse data collection apparatuses and preprocessing methodologies employed by manufacturers. This diversity results in varied data formats, terminological conventions, data fields, and signal recording strategies. Recognizing the challenge this poses, we introduced a unified data representation BatteryData that encompasses both meta information and charge/discharge cycles.
  • Meta information specifications. Battery attributes such as the anode, cathode, and electrolyte materials are recorded as meta information fields. Parameters such as nominal capacity, depth of charge/discharge, and the operating thresholds for voltage and current are also summarized as cell-level attributes.
  • Charge/Discharge cycles. BatteryData organizes the time series records as a list of cycles, detailing charge/discharge protocols and capacity, voltage, current, time, temperature, and internal resistance records.
A comprehensive outline of this BatteryData is presented in Appendix  D.2.
Our commitment to this unified data representation not only facilitates efficient data management from diverse sources but also supports insightful comparisons of battery performance and highlights degradation nuances as battery ages. It creates a conducive environment for deploying machine learning and sophisticated data analysis strategies, propelling advancements in battery data operation and maintenance. For users, BatteryData allows flexible visualization for battery data. As evident in Figure 2, one can glean insights into capacity degradation trends by aggregating all the cycles of the battery cell. Additionally, users can delve into cycle-level signals such as the trajectory of voltage curves of successive cycles or the evolution of Coulombic efficiency for further degradation analysis, as highlighted in Figure 3. Such graphical interpretations provide users with clearer and more intuitive understanding of battery performance and degradation pathways, which in turn benefits the feature and model design.
BatteryML seamlessly supports converting output formats from various battery cycler systems into BatteryData out of the box. Moreover, we offer a suite of automated processing tools to transform existing datasets into the BatteryData format. These dual capabilities ensure BatteryML’s excellent compatibility with a wide range of battery data sources.
@@ 3.2 Feature Engineering
BatteryML boasts an array of degradation features, enabling users to flexibly tailor the data based on distinct experimental needs. These features bifurcate into two categories: within-cycle and between-cycle features.
@@
Within-cycle features. This category encompasses characteristics observed within individual charge/discharge cycles. Examples include
@@
  • QdLinear, which is derived by linear interpolation of the capacity-voltage curve in discharge cycles (Attia et al., 2021).
  • Coulombic efficiency, which is the fraction of charge and discharge capacity within a single cycle, an important indicator of how efficiently a battery can release its stored energy.
@@
  • Internal resistance, which can be calculated by the voltage and current signals to measure the opposition to the flow of electric current within the battery.
@@
Between-cycle features. These features captures the degradation patterns of any battery cell on a higher level, usually across multiple cycles. Examples include
  • Variance of the difference of QdLinear curves, an intuitive yet effective feature that indicates battery degradation speed (Severson et al., 2019).
  • Capacity decay dynamics, the slope of capacity decay curve fitted in early cycles.
  • Average charging time, which reflects the irreversible structural changes within the battery, such as lithium plating and the growth of the SEI layer.
  • Temperature dynamics, which indicates the intensity of the electrochemical reactions occurring within the battery.
  • Minimal internal resistance, reflecting the upper bound of battery health state.
Both the within- and between-cycle features are represented as tensors that are compatible with modern machine learning frameworks. This allows a flexible manipulation of the feature space and a natural combination with deep learning models. We provide a detailed introduction of BatteryML feature module in Appendix D.3. Through these feature extraction methodologies, BatteryML empowers users to freely design and combine features, significantly simplifying the process of reproducing existing work and testing new features.
3.3 Automatic Label Annotation
BatteryML supports automatic label annotation for supervised battery degradation modeling. For different degradation modeling tasks, BatteryML calculates the label as tensors according to the definition. Here we briefly introduce the most important degradation modeling tasks.
As batteries undergo continuous cycles, their capacity inevitably declines due to aging, resulting in slight variances in performance with each use. From the moment of manufacture until a battery is fully aged and retired, a direct and crucial question is to measure the loss of the battery’s maximum available capacity in any given cycle relative to its capacity when new, namely, the cycle’s State of Health (SOH). Specifically, if we denote the battery’s nominal capacity as $C_{\mathrm{nom}}$ and the discharge capacity in the current cycle as $C_{\mathrm{full}}$, the SOH is defined as the percentage of these two quantities
$\mathrm{SOH}=\frac{C_{\mathrm{full}}}{C_{\mathrm{nom}}}\times 100\%.$ (1)
Here the $C_{\mathrm{full}}$ in a strict definition of SOH refers to the capacity measured under the same discharge protocol as the nominal capacity. Note that in practice, this requires additional cycles under constrained temperature and current conditions and is usually infeasible due to significant human labor costs.
This requires accurately estimating the ratio of the capacity already discharged in the current cycle to the original total capacity, that is, the State of Charge (SOC). Essentially, when the depth of discharge is 100%, and the Battery Management System (BMS) has recorded the discharged capacity, SOC is equivalent to SOH. Estimating SOC becomes more challenging when the depth of discharge is less than 100% (for example, when discharging begins again after a brief charge). Specifically, denoting the remaining capacity of the battery as $C_{\mathrm{curr}}$, the definition of SOC is
$\mathrm{SOC}=\frac{C_{\mathrm{curr}}}{C_{\mathrm{full}}}\times 100\%.$ (2)
SOH requires a prediction for each cycle, whereas SOC demands a prediction at every moment during charging and discharging phases. Both tasks requires online prediction, with SOC estimation imposing higher demands on the real-time performance of the model. Another offline task, early battery life prediction, involves conducting a limited number of charge-discharge tests to predict the full lifespan of the battery—typically until its capacity falls below 80% of the nominal capacity (Li et al., 2019)—before significant degradation occurs. This task can greatly shorten the cycling test duration for batteries, playing a crucial role in downstream battery optimization tasks such as the selection of battery materials, optimization of charging strategies, and control of operating temperatures.
BatteryML automates the task of label annotation through sequential traversal of each cycle’s charge/discharge stages, significantly reducing the reliance of model design on battery domain knowledge and allowing machine learning experts to focus on developing superior models.
3.4 Model Development
Following  Attia et al. (2021), BatteryML incorporates multiple off-the-shelf baselines for battery lifetime predictions, including linear models, tree-based models and neural networks. Domain-Enhanced methods like the ’Variance’, ’Discharge’, and ’Full’ models from the  Severson et al. (2019) are implemented using handcrafted features and linear models. BatteryML also includes statistical models such as Ridge regression Hoerl & Kennard (2000), Principal Component Regression(PCR) Tipping & Bishop (1999), Partial Least Squares Regression(PLSR) Geladi & Kowalski (1986), Gaussian process Williams & Rasmussen (2006), XGBoost Chen & Guestrin (2016), Random forest  Breiman (2001) models. For high-performance needs, we offer neural network models like Multi-Layer Perceptron (MLP) Haykin (1994), Convolutional Neural Networks (CNN) Krizhevsky et al. (2012), Long Short-Term Memory networks (LSTM) Hochreiter & Schmidhuber (1997). Additionally, we introduce Transformer Vaswani et al. (2017), a groundbreaking architecture in language and vision domains  (Brown et al., 2020; Dosovitskiy et al., 2021), as a new neural baseline. These implementations utilize scikit-learn Pedregosa et al. (2011) for all statistical models, barring XGBoost, and PyTorch Paszke et al. (2019) for neural networks. We re-train each model with 10 random seeds and report averaged results to eliminate the effect of random initializaiton.
To cater to the diversity of experimental contexts, users have the liberty to tweak and customize these models as per their needs. A deeper dive into model intricacies can be found in AppendixD.5.